# PLUTO-4: Frontier Pathology Foundation Models

### v2

@@ Abstract
Foundation models trained on large-scale pathology image corpora have demonstrated strong transfer capabilities across diverse histopathology tasks. Building on this progress, we introduce PLUTO-4, our next generation of pathology foundation models that extend the Pathology-Universal Transformer (PLUTO) to frontier scale. We share two complementary Vision Transformer architectures in the PLUTO-4 family: a compact and efficient PLUTO-4S model optimized for multi-scale deployment using a FlexiViT setup with 2D-RoPE embeddings, and a frontier-scale PLUTO-4G model trained with a single patch size to maximize representation capacity and stability. Both models are pretrained using a self-supervised objective derived from DINOv2 on a large multi-institutional corpus containing 551,164 WSIs from 137,144 patients across over 50 institutions, spanning over 60 disease types and over 100 stains. Comprehensive evaluation across public and internal benchmarks demonstrates that PLUTO-4 achieves state-of-the-art performance on tasks requiring varying spatial and biological context, including patch-level classification, segmentation, and slide-level diagnosis. The compact PLUTO-4S provides high-throughput and robust performance for practical deployment, while PLUTO-4G establishes new performance frontiers across multiple pathology benchmarks, including 11% improvement in dermatopathology diagnosis. These diverse improvements underscore PLUTO-4’s potential to transform real-world applications as a backbone for translational research and diagnostic use cases.
@@ 1 Introduction
Patch-level Classification; EVA
@@
  • PLUTO-4G, a frontier-scale model trained with a single patch-token size and positional embedding scheme, designed to maximize representation capacity and model performance.
@@
Through extensive benchmarking in patch, slide and segmentation-level tasks, PLUTO-4G shows state-of-the-art performance in multiple public benchmarks, while PLUTO-4S provides a strong and computationally efficient alternative suitable for real-world deployment. Together, these models advance the goal of creating general-purpose pathology foundation models capable of supporting a broad range of diagnostic and research applications.
@@ 3.1 Training Architecture Design
For the frontier-scale model, PLUTO-4G, we observed that larger embeddings inherently possess sufficient representational capacity, and the difference between various patch-token-sizes diminishes as model scale increases. Empirically, we saw the difference in downstream performance between patch-sizes 8, 16 reduce across tasks as model size grows, while the computational cost and memory scales quadratically when patch-token size halves (or sequence length doubles). An example is seen in Figure 2, where we see ViT-G-14 is almost 3.5X faster than ViT-G-8. Therefore, we train PLUTO-4G using a single patch size (14), a choice consistent with other FMs and significantly reduces our training cost while improving throughput.
@@ 4.1 Evaluation Methodology
We evaluate the PLUTO-4 models on a broad set of public and proprietary benchmarks spanning multiple biological scales, including patch-level classification and regression, segmentation, and slide-level prediction. The evaluations were selected to assess the generality of the learned representations across tasks requiring different contextual granularity, ranging from subcellular detail to whole-slide inference.
@@ 4.2 Evaluation Datasets and Reporting Protocols
We evaluate PLUTO-4 models across four broad task categories representing different biological and spatial scales in pathology: patch-level classification, segmentation, spatial transcriptomics, and slide-level prediction. For all benchmarks, we follow standard linear-probing or ABMIL protocols as implemented in EVA (kaiko.ai et al., 2024) and HEST (Jaume et al., 2024).
Patch-level Classification.
@@ 4.3 Benchmark Results
4.3.1 Patch-Level Classification
PLUTO-4G demonstrates strong generalization on a diverse set of patch-level classification tasks (Table 1). On MHIST, PCAM-test and BACH, PLUTO-4G achieves the highest balanced accuracy outperforming other comparable models like Virchow-2, Atlas, H-Optimus-0 and H-Optimus-1. For fine-grained glandular structure classification (Gleason Arvaniti), PLUTO-4G attains the highest accuracy, indicating its ability to model large-scale context. Across general-purpose benchmarks such as BreakHIS and CRC, PLUTO-4G achieves top-tier accuracy, establishing a new performance standard for frozen representation evaluation.
@@ 4.3.4 Slide-Level Prediction
Table 3: Performance results for PLUTO-4: CLS vs. CLS + Mean(Patch-Token) across patch sizes and models. ViT-S models trained with FlexiViT achieve strong performance while allowing configuration of the best patch-token size [8 / 16] for the task. Patch-Token size 8 significantly improves 4S performance on tasks needing fine-grained context.
@@
Patch-level Classification
@@ 4.4 Comparison of PLUTO Models
We next analyze the impact of patch size and embedding type across PLUTO variants (Table 3). This comparison highlights differences between the smaller PLUTO-4S configurations (trained with multiple patch sizes using FlexiViT) and the larger PLUTO-4G model (trained with a single patch size of 14).
@@
Within the ViT-S family, performance differences across patch sizes are modest but consistent. Models trained with smaller patch tokens (ps = 8) slightly outperform ps = 16 on fine-grained, texture-rich datasets such as MHIST, BreakHIS, and BACH, with both CLS, CLS+Mean(Patch-Token) settings, reflecting improved sensitivity to local morphological cues. Conversely, ps = 16 performs on par or better for larger-context tasks such as Gleason, CRC and PCAM. The gap however increases in nuclear segmentation tasks CoNSep and MoNuSAC, where smaller patch-sizes can better capture granular local features with PLUTO-4S-8 performing comparably or better than many larger external FMs like Virchow-2, UNI-2H and closing the gap with PLUTO-4G. These results confirm that flexible patch-token training effectively balances global and local context while providing flexibility to select the best patch-token size for modelling depending on the task.
@@
The larger PLUTO-4G model delivers substantial gains across all task categories. It surpasses both ViT-S configurations on all patch-level benchmarks and achieves the best results on segmentation, with Dice/AJI scores of 70.4/65.0, indicating stronger spatial consistency in feature maps. This scaling trend mirrors the improvements observed in other large-scale vision transformers, where increased capacity enhances representational generality across tasks.
@@
Overall, the addition of patch-level context yields modest but measurable benefits for classification tasks, particularly in smaller ViT-S models, while larger ViT-G models appear less sensitive to the choice of embedding aggregation strategy.
@@ 5 Discussion
The PLUTO-4 series demonstrates the next step in scaling and adapting foundation models for digital pathology. Our results establish PLUTO-4G as a new performance standard across diverse pathology benchmarks, achieving state-of-the-art results in patch-level, segmentation, and slide-level prediction tasks. Trained at scale with a large, diverse corpus of histopathology data, PLUTO-4G exhibits strong generalization across datasets and clinical contexts, capturing both cellular- and tissue-level morphology within a unified representation. In parallel, PLUTO-4S provides an efficient and versatile alternative—offering competitive performance with substantially reduced compute requirements and configurable patch-token sizes that make it well suited for a wide range of downstream problems and deployment settings.

### v3

@@ Abstract
Foundation models trained on large-scale pathology image corpora have demonstrated strong transfer capabilities across diverse histopathology tasks. Building on this progress, we introduce PLUTO-4, our next generation of pathology foundation models that extend the Pathology-Universal Transformer (PLUTO) to frontier scale. We share two complementary Vision Transformer architectures in the PLUTO-4 family: a compact and efficient PLUTO-4S model optimized for multi-scale deployment using a FlexiViT setup with 2D-RoPE embeddings, and a frontier-scale PLUTO-4G model trained with a single patch size to maximize representation capacity and stability. Both models are pretrained using a self-supervised objective derived from DINOv2 on a large multi-institutional corpus containing 551,164 WSIs from 137,144 patients across over 50 institutions, spanning over 60 disease types and over 100 stains. Comprehensive evaluation across public and internal benchmarks demonstrates that PLUTO-4 achieves state-of-the-art performance on tasks requiring varying spatial and biological context, including tile classification, segmentation, and slide-level diagnosis. The compact PLUTO-4S provides high-throughput and robust performance for practical deployment, while PLUTO-4G establishes new performance frontiers across multiple pathology benchmarks, including 11% improvement in dermatopathology diagnosis. These diverse improvements underscore PLUTO-4’s potential to transform real-world applications as a backbone for translational research and diagnostic use cases.
@@ 1 Introduction
Tile-Level Classification; EVA
@@
  • PLUTO-4G, a frontier-scale model trained with a single patch size (also referred as patch-token size) and positional embedding scheme, designed to maximize representation capacity and model performance.
@@
Through extensive benchmarking in tile, slide and segmentation-level tasks, PLUTO-4G shows state-of-the-art performance in multiple public benchmarks, while PLUTO-4S provides a strong and computationally efficient alternative suitable for real-world deployment. Together, these models advance the goal of creating general-purpose pathology foundation models capable of supporting a broad range of diagnostic and research applications.
@@ 3.1 Training Architecture Design
For the frontier-scale model, PLUTO-4G, we observed that larger embeddings inherently possess sufficient representational capacity, and the difference between various patch-token sizes diminishes as model scale increases. Empirically, we saw the difference in downstream performance between patch-token sizes 8, 16 reduce across tasks as model size grows, while the computational cost and memory scales quadratically when patch-token size halves (or sequence length doubles). An example is seen in Figure 2, where we see ViT-G-14 is almost 3.5X faster than ViT-G-8. Therefore, we train PLUTO-4G using a single patch-token size (14), a choice consistent with other FMs and significantly reduces our training cost while improving throughput.
@@ 4.1 Evaluation Methodology
We evaluate the PLUTO-4 models on a broad set of public and proprietary benchmarks spanning multiple biological scales, including Tile-Level classification and regression, segmentation, and slide-level prediction. The evaluations were selected to assess the generality of the learned representations across tasks requiring different contextual granularity, ranging from subcellular detail to whole-slide inference.
@@ 4.2 Evaluation Datasets and Reporting Protocols
We evaluate PLUTO-4 models across four broad task categories representing different biological and spatial scales in pathology: Tile-Level classification, segmentation, spatial transcriptomics, and slide-level prediction. For all benchmarks, we follow standard linear-probing or ABMIL protocols as implemented in EVA (kaiko.ai et al., 2024) and HEST (Jaume et al., 2024).
Tile-Level Classification.
@@ 4.3 Benchmark Results
4.3.1 Tile-Level Classification
PLUTO-4G demonstrates strong generalization on a diverse set of Tile-Level classification tasks (Table 1). On MHIST, PCAM-test and BACH, PLUTO-4G achieves the highest balanced accuracy outperforming other comparable models like Virchow-2, Atlas, H-Optimus-0 and H-Optimus-1. For fine-grained glandular structure classification (Gleason Arvaniti), PLUTO-4G attains the highest accuracy, indicating its ability to model large-scale context. Across general-purpose benchmarks such as BreakHIS and CRC, PLUTO-4G achieves top-tier accuracy, establishing a new performance standard for frozen representation evaluation.
@@ 4.3.4 Slide-Level Prediction
Table 3: Performance results for PLUTO-4: CLS vs. CLS + Mean(Patch-Token) across patch-token sizes and models. ViT-S models trained with FlexiViT achieve strong performance while allowing configuration of the best patch-token size [8 / 16] for the task. Patch-Token size 8 significantly improves 4S performance on tasks needing fine-grained context, while patch-token size of 16 can perform on-par or better than patch-token size 8 on many coarse grained tasks while being 4X faster.
@@
Tile-Level Classification
@@ 4.4 Comparison of PLUTO Models
We next analyze the impact of patch-token size (also referred as patch size) and embedding type across PLUTO variants (Table 3). This comparison highlights differences between the smaller PLUTO-4S configurations (trained with multiple patch-token sizes using FlexiViT) and the larger PLUTO-4G model (trained with a single patch-token size of 14).
@@
Within the ViT-S family, performance differences across patch-token sizes are modest but consistent. Models trained with smaller patch-tokens (ps = 8) slightly outperform ps = 16 on fine-grained, texture-rich datasets such as MHIST, BreakHIS, and BACH, with both CLS, CLS+Mean(Patch-Token) settings, reflecting improved sensitivity to local morphological cues. Conversely, ps = 16 performs on par or better for larger-context tasks such as Gleason, CRC and PCAM. The gap however increases in nuclear segmentation tasks CoNSep and MoNuSAC, where smaller patch-sizes can better capture granular local features with PLUTO-4S-8 performing comparably or better than many larger external FMs like Virchow-2, UNI-2H and closing the gap with PLUTO-4G. These results confirm that flexible patch-token training effectively balances global and local context while providing flexibility to select the best patch-token size for modelling depending on the task.
@@
The larger PLUTO-4G model delivers substantial gains across all task categories. It surpasses both ViT-S configurations on all Tile-Level benchmarks and achieves the best results on segmentation, with Dice/AJI scores of 70.4/65.0, indicating stronger spatial consistency in feature maps. This scaling trend mirrors the improvements observed in other large-scale vision transformers, where increased capacity enhances representational generality across tasks.
@@
Overall, the addition of Tile-Level context yields modest but measurable benefits for classification tasks, particularly in smaller ViT-S models, while larger ViT-G models appear less sensitive to the choice of embedding aggregation strategy.
@@ 5 Discussion
The PLUTO-4 series demonstrates the next step in scaling and adapting foundation models for digital pathology. Our results establish PLUTO-4G as a new performance standard across diverse pathology benchmarks, achieving state-of-the-art results in Tile-Level, segmentation, and slide-level prediction tasks. Trained at scale with a large, diverse corpus of histopathology data, PLUTO-4G exhibits strong generalization across datasets and clinical contexts, capturing both cellular- and tissue-level morphology within a unified representation. In parallel, PLUTO-4S provides an efficient and versatile alternative—offering competitive performance with substantially reduced compute requirements and configurable patch-token sizes that make it well suited for a wide range of downstream problems and deployment settings.